# The Nesterov–Spokoiny Acceleration Achieves Strict o(1/k2) Convergence

### v4

@@ Abstract
We study a variant of an accelerated gradient algorithm of Nesterov and Spokoiny. We call this algorithm the Nesterov–Spokoiny Acceleration (NSA). The NSA algorithm simultaneously satisfies the following property: For a smooth convex objective $f\in\mathscr{F}_{L}^{\infty,1}(\mathbb{R}^{n})$, the sequence $\{\mathbf{x}_{k}\}_{k\in\mathbb{N}}$ governed by NSA satisfies $\limsup\limits_{k\to\infty}k^{2}(f(\mathbf{x}_{k})-f^{*})=0$, where $f^{*}>-\infty$ is the minimum of $f$.
@@ 1 Introduction
A central problem in optimization and related fields is solving
@@
where $f$ is a real-valued convex function bounded from below. An efficient solver for (1) has important real-world implications, since (1) captures the core challenge of many important problems, ranging from systems control Lewis et al., (2012) to neural network training Goodfellow et al., (2016). In neural network training, gradients can be calculated using back-propagation, while function values are obtained through forward-propagation. The back-propagation step is often significantly more resource-intensive than forward-propagation. As a result, it is often more advantageous to fine-tune large-scale neural networks using zeroth-order optimization methods, which only require function evaluations.
@@ 1.1 Zeroth-order Optimization
Recently, the rising demand for fine-tuning large language models has brought zeroth-order optimization back into the spotlight (Zhang et al., (2025), and references therein). This resurgence is largely due to the memory efficiency of zeroth-order optimization, making it particularly suitable for large-scale deep learning challenges. Although there is a vast body of literature on zeroth-order optimization (refer to Conn et al., (2009) for a contemporary overview), which encompasses techniques including Bayesian optimization Shahriari et al., (2015), stochastic optimization Robbins and Monro, (1951); Kiefer and Wolfowitz, (1952); Nemirovski et al., (2009) and genetic algorithms Srinivas and Patnaik, (1994), the application of zeroth-order optimization in fine-tuning is quite straightforward. We estimate the gradient using zeroth-order information and apply gradient-based methods using the estimated gradient.
@@
Over the years, gradient-based algorithms have become the default solver for (1) in several scenarios, and a tremendous amount of gradient methods have been invented. The history of gradient algorithms dates back to classical times, when Cauchy proposed a method of finding the minimum value of a function via first-order information Lemaréchal, (2012). Nowadays, gradient-based zeroth-order methods are gaining importance.
Throughout the years, many gradient-based zeroth-order optimization methods have been studied (e.g.,Flaxman et al., (2005); Ghadimi and Lan, (2013); Duchi et al., (2015); Nesterov and Spokoiny, (2017); Wang et al., (2018); Liu et al., (2018); Ji et al., (2019); Balasubramanian and Ghadimi, (2021); Yue et al., (2023); Wang et al., (2023)); See Liu et al., (2020) for an exposition. For example, Flaxman et al., (2005) studied the stochastic gradient estimator using a single-point function evaluation for the purpose of bandit learning. Duchi et al., (2015) studied stabilization of the stochastic gradient estimator via two-points (or multi-points) evaluations. Nesterov and Spokoiny, (2017); Balasubramanian and Ghadimi, (2021) studied gradient/Hessian estimators using Gaussian smoothing, and investigated downstream optimization methods using the estimated gradient. In particular, Nesterov and Spokoiny, (2017) introduced the first accelerated zeroth-order method that achieves an $O(k^{-2})$ convergence rate in expectation.
@@ 1.2 Accelerated Gradient Methods
Based on a common physics phenomenon and the original Gradient Algorithm, Polyak’s Heavy-ball method was invented Polyak, (1964). This algorithm is perhaps the first method that directly aims to accelerate the convergence of gradient search. One of the most surprising works in this field is the Nesterov’s Accelerated Gradient (NAG) method Nesterov, (1983), where accelerated gradient methods were invented to speed up the convergence for smooth convex programs. For such problems, NAG achieves a $O(k^{-2})$ convergence rate, without incurring computations of higher-order expense. When the iteration counter $k$ is not large, the convergence rate of NAG is proved to be optimal among methods where the next iteration can be written as a linear combination of previous gradients and the starting point Nesterov, (2018).
@@
Since Nesterov’s original acceleration method Nesterov, (1983), several accelerated gradient methods have been invented, including the one that inspires our work Nesterov, (2011); Nesterov and Spokoiny, (2017). Also, Flammarion and Bach, (2015) builds a connection between average gradient and acceleration. The nature of the acceleration phenomenon has also been investigated from a continuous-time perspective. In particular, Su et al., (2016) pointed out the relation between the a second-order ordinary differential equation and NAG. This discovery allows for a better understanding of Nesterov’s scheme. Afterwards, a family of schemes with $O(1/k^{2})$ convergence rates has been obtained. Shi et al., (2022) finds a finer correspondence between discrete-time iteration and continuous-time trajectory and proposes a gradient-correction scheme of NAG. Chen et al., (2022) introduced an implicit-velocity scheme of NAG and investigates on the convergence rate of gradient norms. Please See Section 4.1.1 for more discussions on the continuous-time perspective.
@@
Proximal operations have been widely employed in acceleration methods; See e.g., Parikh et al., (2014) for an exposition. The proximal operator can handle possible nonsmoothness in the objective. Roughly speaking, the proximal operator finds the next iteration point by minimizing an approximation of the objective (or part of the objective) plus a regularizer term. In a broad sense, the mirror map in mirror descent Nemirovsky and Yudin, (1983) is also a type of proximal operator. In discussing related works, mirror descent algorithms are considered as proximal methods.
@@
Over the years, various accelerated gradient/proximal methods have been invented. Nesterov, (2005) introduced a proximal method that achieves $O(k^{-2})$ convergence rate for convex nonsmooth programs. Beck and Teboulle, (2009) popularized accelerated proximal methods via the Fast Iterative Shrinkage-Thresholding Algorithm (FISTA) algorithm, which achieves a better rate of convergence for linear inverse problems, while preserving the computational simplicity of ISTA Combettes and Wajs, (2005); Daubechies et al., (2004). FISTA has also been analyzed in Chambolle and Dossal, (2015). Flammarion and Bach, (2015) shows that average gradient methods can also achieve acceleration. (Attouch et al., , 2022) introduced the Inertial Proximal Algorithm with Hessian Damping (IPAHD), an accelerated first-order method that utilizes the physics of Hessian damping. Motivated by the complementary performances of mirror descent and gradient descent, Allen-Zhu and Orecchia, (2014); Diakonikolas and Orecchia, (2019) couple these two celebrated algorithms and present a new scheme to reconstruct NAG on the basis of other algorithms. Krichene et al., 2015a builds a continuous-time analysis of the accelerated mirror descent method, whose discretized version also converges at rate $O(\frac{1}{k^{2}})$. Bubeck et al., (2015) introduced a modification of NAG with a geometric interpretation and a proof of the same convergence rate as NAG. Notably, Attouch and Peypouquet, (2016) proved that the forward-backward method of Nesterov Attouch et al., (2014) converges at rate $o(k^{-2})$.
@@
However, one notable drawback of all Nesterov-type acceleration methods, when compared to gradient descent, is that the function value does not consistently decrease. This phenomenon arises from the momentum introduced during the gradient step. In this paper, we present a method that combines the benefits of both acceleration momentum and descent properties. Our approach is a variant of the acceleration method developed by Nesterov and Spokoiny, and achieves a convergence rate of $o(1/k^{2})$. Also, we prove a $o(k^{-2})$ convergence rate of a zeroth-order implementation of this algorithm.
@@ 1.3 Our Contributions
In this paper, we study an acceleration gradient method called Nesterov–Spokoiny Acceleration (NSA), which is a variation of an acceleration method of Nesterov and Spokoiny Nesterov, (2011); Nesterov and Spokoiny, (2017). Our method harnesses the advantages of both accelerated gradient methods and traditional gradient descent – it leverages an acceleration momentum, and ensures the decrease of function values. Theoretically, the NSA algorithm satisfies the following convergence rate properties:
@@
    where $f^{*}>-\infty$ is the minimum of the smooth convex function $f$.
@@
    The sequence $\{\mathbf{x}_{k}\}_{k\in\mathbb{N}}$ governed by NSA satisfies
    $\liminf\limits_{k\to\infty}k^{2}\log k\log\log k(f(\mathbf{x}_{k})-f^{*})=0.$
  3. 3.
@@
Furthermore, we demonstrate that the NSA algorithm allows for a zeroth-order implementation that attains the same convergence rates. We also present a continuous-time analysis of the NSA algorithm, and establishes a convergence rate for the high-resolution ODE Shi et al., (2022) in the convex setting (which is not necessarily strongly convex). This closed a gap identified in Shi et al., (2022). For further details, please refer to Proposition 2 and Remark 2. Comparison of NSA with state-of-the-art prior arts is summarized in Table 1.
Extensions of the NSA algorithm is also studied. In Section 3, we study a variation of the NSA algorithm that handles inexact gradient. In Section 3, we show that a zeroth-order variant of NSA also achieves $o(1/k^{2})$ convergence. In Section 4, we study a variation of the NSA algorithm that solves convex nonsmooth programs. A continuous-time analysis is also provided. Some more literature from the continuous-time perspective are discussed in Section 4.1.
[…]
Nesterov, (1983) $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$
Beck and Teboulle, (2009) $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$ N/A
Allen-Zhu and Orecchia, (2014) $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$ N/A
Flammarion and Bach, (2015) $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$
Krichene et al., 2015a $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$
Krichene et al., 2015b $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$ N/A
Attouch and Peypouquet, (2016) $o(\frac{1}{k^{2}})$ $o(\frac{1}{k^{2}\log k})$ N/A
Nesterov and Spokoiny, (2017) $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$
Diakonikolas and Orecchia, (2019) $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$
Attouch, Hedy et al., (2019) $o(\frac{1}{k^{2}})$ $o(\frac{1}{k^{2}})$ N/A
Attouch et al., (2022) $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$ N/A
Shi et al., (2022) $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{3}})$
Chen et al., (2022) $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$ $o(\frac{1}{k^{3}})$
Su et al., (2016) $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{2}})$ N/A
This work $o(\frac{1}{k^{2}})$ $o(\frac{1}{k^{2}\log k\log\log k})$ $o(\frac{1}{k^{3}\log k\log\log k})$
Table 1: Comparison with state-of-the-art acceleration methods for convex programs, listed in chronological order. The third column for proximal methods that targets at nonsmooth objectives are filled in with “N/A”. Many of the existing works have other virtues. For example, some find fine physical formulations of the acceleration phenomenon Su et al., (2016); Shi et al., (2022); some build prime-dual interpretation of the acceleration phenomenon Allen-Zhu and Orecchia, (2014); Diakonikolas and Orecchia, (2019); some handles nonsmooth objectives and composite optimization problems Beck and Teboulle, (2009); Attouch and Peypouquet, (2016); Attouch, Hedy et al., (2019). Su et al., (2016) enjoys an $O(k^{-3})$ convergence rate for strongly convex problems without any knowledge about the strong-convexity parameter. Yet in terms of convergence rate among methods that admits zeroth-order implementation, the NSA algorithm studied in this paper is the best. Also, NSA simultaneously achieves $o(k^{-2})$ convergence rate for the function value and $o(k^{-3})$ convergence rate for the gradient norm.
@@ 2 The Nesterov–Spokoiny Acceleration
In the work of Nesterov and Spokoiny Nesterov and Spokoiny, (2017), an acceleration principle that differs from Nesterov’s classic acceleration method Nesterov, (1983, 2003) is presented. This method involves maintaining two sequences, $\{\mathbf{x}_{k}\}$ and $\{\mathbf{z}_{k}\}$, and utilizes the gradient on the line connecting $\mathbf{x}_{k}$ and $\mathbf{z}_{k}$ to update both trajectories. Nesterov and Spokoiny introduced this form of acceleration specifically in the context of zeroth-order acceleration. To effectively enhance this method of Nesterov and Spokoiny, we integrate the principle of gradient descent into this acceleration method. A similar principle previously appeared in Li et al., (2017). In this paper, we demonstrate that this design leverages the advantages of the descent property and acceleration momentum, resulting in a convergence rate of $o(k^{-2})$. The Nesterov–Spokoiny Acceleration algorithm is described in Algorithm 1.
@@
4:   $\mathbf{x}_{k+1}^{\prime}=\mathbf{y}_{k}-\eta\nabla f(\mathbf{y}_{k})$. /* $\eta$ is the step size. */
5:   $\mathbf{x}_{k+1}^{\prime\prime}=\mathbf{x}_{k}-\eta\nabla f(\mathbf{x}_{k})$.
6:   $\mathbf{x}_{k+1}=\begin{cases}\mathbf{x}_{k+1}^{\prime},&\text{ if }f(\mathbf{x}_{k+1}^{\prime})\leq f(\mathbf{x}_{k+1}^{\prime\prime})\\ \mathbf{x}_{k+1}^{\prime\prime},&\text{ otherwise. }\end{cases}$
7:   $\mathbf{z}_{k+1}=\mathbf{z}_{k}-\frac{\eta}{\alpha_{k}}\nabla f(\mathbf{y}_{k})$.
@@
Figure 1: The figure above illustrates a comparison between Algorithm 1 and the original algorithm by Nesterov and Spokoiny Nesterov, (2017), which shows their paths in one iteration from the same starting point. The curves represent the contour lines. The points that are underlined are the new iteration points of Algorithm 1, while the rest new points are from the original algorithm by Nesterov and Spokoiny Nesterov, (2017). The red and green dashed lines represent the function values of the next step for $x_{t}$, respectively.
Convergence guarantee for smooth convex objectives $\mathscr{F}_{L}^{1,1}(\mathbb{R}^{n})$ Nesterov, (2003, 2018) is in Theorem 1. For completeness, the definition of $\mathscr{F}_{L}^{1,1}(\mathbb{R}^{n})$ is included in the Appendix.
@@
Instate all notations in Algorithm 1. Consider an objective function $f\in\mathscr{F}_{L}^{1,1}(\mathbb{R}^{n})$. Let $\eta\in(0,\frac{2}{3L}]$ and let $p\geq 3$. Let $\alpha_{k}=\frac{p}{k+p}$ for each $k\in\mathbb{N}$.
@@
Let $f\in\mathscr{F}_{L}^{1,1}(\mathbb{R}^{n})$ and let $\eta\in(0,\frac{2}{3L}]$. Let $\mathbf{x}_{k},\mathbf{y}_{k}$ be governed by the NSA algorithm. Then for all $t=0,1,2,\cdots$, it holds that
$\delta_{k+1}\leq\delta_{k}\quad\text{ and }\quad f(\mathbf{x}_{k+1})\leq f(\mathbf{y}_{k})-\frac{2\eta}{3}\|\nabla f(\mathbf{y}_{k})\|^{2},$
where $\delta_{k}:=f(\mathbf{x}_{k})-f^{*}$ is introduced to avoid notation clutter.
@@
By (2) we know
$f(\mathbf{x}_{k+1})\leq f(\mathbf{x}_{k+1}^{\prime})\leq f(\mathbf{y}_{k})-\frac{2\eta}{3}\|\nabla f(\mathbf{y}_{k})\|^{2}.$
@@
Proof of Theorem 1.
@@
where $\gamma_{k+1}=\frac{\alpha_{k}^{2}}{\eta}$ for all $t=0,1,\cdots$. Thus for any $t\geq 1$, we have

### v5

@@ Abstract
This paper studies the Nesterov-Spokoiny Acceleration (NSA), a variant of the accelerated gradient method by Nesterov and Spokoiny. For smooth convex optimization, NSA achieves a strict $o(1/k^{2})$ convergence rate in function value and an $o(1/(k^{3}\log k))$ rate in squared gradient norm, while ensuring monotonic descent of the objective. We further study a zeroth-order version of NSA that handles inexact gradients, and extends NSA to composite optimization problems, in each case establishing $o(1/k^{2})$ convergence in function value. A continuous-time analysis reveals connections to high-resolution ODEs known to underlie acceleration phenomena.
keywords:
Continuous Optimization, Convex Optimization
journal: European Journal of Operational Research
[…]
[fdu1]organization=School of Mathematics, Fudan University, city=Shanghai, country=China
[…]
[fdu2]organization=Shanghai Center for Mathematica Sciences, Fudan University, city=Shanghai, country=China
@@ 1 Introduction
A central challenge in optimization is efficiently solving the problem
@@
where $f$ is a convex function bounded from below. The pursuit of robust solvers for this problem is driven by its profound real-world utility, as it encapsulates the core computational task in fields ranging from machine learning—where it facilitates model training via empirical risk minimization [jordan2015machine]—to systems control [lewis2012optimal], signal processing [orfanidis1995introduction], and operations research [pennanen2012introduction]. Consequently, advancements in convex optimization directly translate to enhanced performance and scalability across much of the modern technological landscape.
Given the widespread applicability of (1), developing efficient algorithms for its solution is of paramount importance. The most straightforward approach, gradient descent, often suffers from a slow convergence rate, particularly for ill-conditioned problems. Second-order methods, like Newton’s method [e.g., polyak2007newton, and references therein], address this by using the Hessian to achieve superlinear convergence, but at a prohibitive computational cost for large-scale problems and with only local convergence guarantees.
The quest to replicate this acceleration has therefore driven the development of advanced first-order methods. This pursuit is highly relevant to the rise of gradient-based zeroth-order optimization, which is particularly effective for fine-tuning large neural networks. In this context, the cost of computing gradients through back-propagation heavily outweighs the cost of a simple function evaluation via forward-propagation [malladi2023finetuning]. Therefore, advances in efficient first-order methods potentially inform and inspire the development of better zeroth-order optimizers.
@@ 1.1 Zeroth-order Optimization
Recently, the rising demand for fine-tuning large language models has brought zeroth-order optimization back into the spotlight [e.g., malladi2023finetuning, 10.5555/3692070.3694514]. This resurgence is largely due to the memory efficiency of zeroth-order optimization, making it particularly suitable for large-scale deep learning challenges.
Although there is a vast body of literature on zeroth-order optimization (refer to [conn2009introduction] for a contemporary overview), which encompasses techniques including Bayesian optimization [shahriari2015taking], stochastic optimization [robbins1951stochastic, kiefer1952stochastic, nemirovski2009robust] and genetic algorithms [srinivas1994genetic], the application of zeroth-order optimization in fine-tuning is quite straightforward. We estimate the gradient using zeroth-order information and apply gradient-based methods using the estimated gradient.
@@
Over the years, gradient-based algorithms have become the default solver for (1) in several scenarios, and a tremendous amount of gradient methods have been invented. The history of gradient algorithms dates back to classical times, when Cauchy proposed a method of finding the minimum value of a function via first-order information [lemarechal2012cauchy]. Nowadays, gradient-based zeroth-order methods are gaining importance.
Throughout the years, many gradient-based zeroth-order optimization methods have been studied (e.g.,[flaxman2005online, ghadimi2013stochastic, duchi2015optimal, nesterov2017random, wang2018stochastic, liu2018zeroth, ji2019improved, balasubramanian2021zeroth, yue2023zeroth, Wang2023]); See [liu2020primer] for an exposition. For example, [flaxman2005online] studied the stochastic gradient estimator using a single-point function evaluation for the purpose of bandit learning. [duchi2015optimal] studied stabilization of the stochastic gradient estimator via two-points (or multi-points) evaluations. nesterov2017random, balasubramanian2021zeroth studied gradient/Hessian estimators using Gaussian smoothing, and investigated downstream optimization methods using the estimated gradient. In particular, [nesterov2017random] introduced the first accelerated zeroth-order method that achieves an $O(k^{-2})$ convergence rate in expectation.
@@ 1.2 Accelerated Gradient Methods
Based on a common physics phenomenon and the original Gradient Algorithm, Polyak’s Heavy-ball method was invented [1964Some]. This algorithm is perhaps the first method that directly aims to accelerate the convergence of gradient search. One of the most surprising works in this field is the Nesterov’s Accelerated Gradient (NAG) method [nesterov1983method], where accelerated gradient methods were invented to speed up the convergence for smooth convex programs. For such problems, NAG achieves a $O(k^{-2})$ convergence rate, without incurring computations of higher-order expense. Furthermore, for bounded iteration counts $k$, the convergence rate of NAG is provably optimal among all methods whose iterates are constrained to the span of previous gradients and the initial point [nest2018].
@@
Since Nesterov’s original acceleration method [nesterov1983method], several accelerated gradient methods have been invented, including the one that inspires our work [nesterov2011random, nesterov2017random]. Also, [Flammarion2015FromAT] builds a connection between average gradient and acceleration. Beyond these discrete-time analyses, the fundamental nature of the acceleration phenomenon has been fruitfully investigated from a continuous-time perspective. In particular, [JMLR:v17:15-084] pointed out the relation between the a second-order ordinary differential equation and NAG. This discovery allows for a better understanding of Nesterov’s scheme. Afterwards, a family of schemes with $O(1/k^{2})$ convergence rates has been obtained. [2018Understanding] finds a finer correspondence between discrete-time iteration and continuous-time trajectory and proposes a gradient-correction scheme of NAG. [chen2022gradient] introduced an implicit-velocity scheme of NAG and investigates on the convergence rate of gradient norms. Please See Section 4.1.1 for more discussions on the continuous-time perspective.
@@
The principles of acceleration have been successfully combined with other powerful optimization tools, most notably the proximal operator; See e.g., [parikh2014proximal] for an exposition. The proximal operator can handle possible nonsmoothness in the objective. Roughly speaking, the proximal operator finds the next iteration point by minimizing an approximation of the objective (or part of the objective) plus a regularizer term. In a broad sense, the mirror map in mirror descent [blair1985problem] is also a type of proximal operator. In discussing related works, mirror descent algorithms are considered as proximal methods.
@@
Over the years, various accelerated gradient/proximal methods have been invented. [nesterov2005smooth] introduced a proximal method that achieves $O(k^{-2})$ convergence rate for convex nonsmooth programs. [2009A] popularized accelerated proximal methods via the Fast Iterative Shrinkage-Thresholding Algorithm (FISTA) algorithm, which achieves a better rate of convergence for linear inverse problems, while preserving the computational simplicity of ISTA [combettes2005signal, daubechies2004iterative]. FISTA has also been analyzed in [chambolle2015convergence]. [Flammarion2015FromAT] shows that average gradient methods can also achieve acceleration. [attouch2022first] introduced the Inertial Proximal Algorithm with Hessian Damping (IPAHD), an accelerated first-order method that utilizes the physics of Hessian damping. Motivated by the complementary performances of mirror descent and gradient descent, [allen2014linear, doi:10.1137/18M1172314] couple these two celebrated algorithms and present a new scheme to reconstruct NAG on the basis of other algorithms. [10.5555/2969442.2969558] builds a continuous-time analysis of the accelerated mirror descent method, whose discretized version also converges at rate $O(k^{-2})$. [bubeck2015geometric] introduced a modification of NAG with a geometric interpretation and a proof of the same convergence rate as NAG. Notably, [doi:10.1137/15M1046095] proved that the forward-backward method of Nesterov [doi:10.1137/130910294] converges at rate $o(k^{-2})$.
@@
However, one notable drawback of all Nesterov-type acceleration methods, when compared to gradient descent, is that the function value does not consistently decrease. This phenomenon arises from the momentum introduced during the gradient step. This paper presents a method that unifies the benefits of acceleration momentum and descent properties. While techniques for ensuring monotonicity in accelerated algorithms exist—such as the monotone APG by [beck2009fast2] and subsequent APG-type methods by [li2015accelerated] which ensure sufficient descent for nonconvex problems while maintaining an $O(k^{-2})$ rate for convex ones—their nonconvex convergence was later analyzed under the KŁ framework by [li2017convergence]. We are the first to demonstrate that such monotone methods simultaneously achieve a $o(k^{-2})$ $\limsup$ convergence in function value and $o(\frac{1}{k^{3}\log k})$ $\liminf$ convergence in gradient norm. Furthermore, we prove that a zeroth-order implementation of our algorithm also enjoys a $o(k^{-2})$ convergence in function value.
@@ 1.3 Our Contributions
In this paper, we study an acceleration gradient method called Nesterov–Spokoiny Acceleration (NSA), which is a variation of an acceleration method of Nesterov and Spokoiny [nesterov2011random, nesterov2017random]. Our method harnesses the advantages of both accelerated gradient methods and traditional gradient descent – it leverages an acceleration momentum, and ensures the decrease of function values. Theoretically, the NSA algorithm satisfies the following convergence rate properties:
@@
    and $\liminf\limits_{k\to\infty}k^{2}\log k\log\log k(f(\mathbf{x}_{k})-f^{*})=0$, where $f^{*}>$-$\infty$ is the minimum of the smooth convex function $f$.
@@
To our knowledge, this work is the first to establish an $o(k^{-2})$ convergence rate for the function value and an $o(\frac{1}{k^{3}\log k})$ rate for the squared gradient norm, all while maintaining strictly monotonic descent of function value.
Also, this work presents a comprehensive study of the NSA algorithm. First, we introduce a zeroth-order variant capable of handling inexact gradients and prove its convergence at an $o(1/k^{2})$ rate. Second, we propose a modification designed for convex nonsmooth problems, and provide a continuous-time analysis of its dynamics. Lastly, we adapt the framework to the composite optimization setting, where it likewise achieves an $o(1/k^{2})$ convergence rate in the function value. A comparison of NSA with state-of-the-art methods is summarized in Table 1.
$\limsup$ converg. rate for func. value $\liminf$ converg. rate for grad. norm sqr. grad. comp. per iter. guaranteed monot. descent? explicitly support prox.?
[nesterov1983method] $O(\frac{1}{k^{2}})$ N/A 1 No N/A
[2009A] $O(\frac{1}{k^{2}})$ N/A 1 No Yes
[allen2014linear] $O(\frac{1}{k^{2}})$ N/A 1 No Yes
[Flammarion2015FromAT] $O(\frac{1}{k^{2}})$ N/A 1 No N/A
[10.5555/2969442.2969558] $O(\frac{1}{k^{2}})$ N/A 1 No Yes
[JMLR:v17:15-084] $O(\frac{1}{k^{2}})$ N/A 1 No Yes
[doi:10.1137/15M1046095] $o(\frac{1}{k^{2}})$ N/A 1 No Yes
[nesterov2017random] $O(\frac{1}{k^{2}})$ N/A 1 No N/A
[doi:10.1137/18M1172314] $O(\frac{1}{k^{2}})$ N/A 1 No Yes
[2019rate] $o(\frac{1}{k^{2}})$ N/A 1 No Yes
[attouch2022first] $O(\frac{1}{k^{2}})$ $O(\frac{1}{k^{3}})$ 1 No Yes
[2018Understanding] $o(\frac{1}{k^{2}})$ $O(\frac{1}{k^{3}})$ 1 No N/A
[chen2022gradient] $O(\frac{1}{k^{2}})$ $o(\frac{1}{k^{3}})$ 1 No N/A
This work Theorem 1, 6 $o(\frac{1}{k^{2}})$ $o(\frac{1}{k^{3}\log k})$ 2 Yes Yes
Table 1: This table provides a chronological comparison of state-of-the-art acceleration methods for convex programs. The columns report key theoretical properties: the $\limsup$ convergence rate for the function value, the $\liminf$ convergence rate for the squared gradient norm, the number of gradient computations per iteration (excluding proximal operations), whether the algorithm ensures a monotonically decreasing objective value, and whether a proximal version has been explicitly analyzed. Entries marked "N/A" indicate that the corresponding result is not known or not applicable. Notably, our work is the first to simultaneously achieve $o(k^{-2})$ $\limsup$ convergence in function value, and $o(\frac{1}{k^{3}\log k})$ $\liminf$ convergence in squared gradient norm.
@@ 2 The Nesterov–Spokoiny Acceleration
In the work of Nesterov and Spokoiny [nesterov2017random], an acceleration principle that differs from Nesterov’s classic acceleration method [nesterov1983method, nest2003] is presented. This method involves maintaining two sequences, $\{\mathbf{x}_{k}\}$ and $\{\mathbf{z}_{k}\}$, and utilizes the gradient on the line connecting $\mathbf{x}_{k}$ and $\mathbf{z}_{k}$ to update both trajectories. Nesterov and Spokoiny introduced this form of acceleration specifically in the context of zeroth-order acceleration. To effectively enhance this method of Nesterov and Spokoiny, we integrate the principle of gradient descent into this acceleration method. A similar principle previously appeared in [pmlr-v70-li17g]. In this paper, we provide a design that merges the descent property with Nesterov’s momentum, achieving a convergence rate of $o(k^{-2})$ while maintaining monotonic descent. The Nesterov–Spokoiny Acceleration algorithm is described in Algorithm 1.
@@
[…]
@@
Figure 1: The figure above illustrates a comparison between Algorithm 1 and the original algorithm by Nesterov and Spokoiny [2017Random], which shows their paths in one iteration from the same starting point. The curves represent the contour lines. The points that are underlined are the new iteration points of Algorithm 1, while the rest new points are from the original algorithm by Nesterov and Spokoiny [2017Random]. The red and green dashed lines represent the function values of the next step for $x_{t}$, respectively.
Convergence guarantee for smooth convex objectives $\mathscr{F}_{L}^{1,1}(\mathbb{R}^{n})$ [nest2003, nest2018] is in Theorem 1. For completeness, the definition of $\mathscr{F}_{L}^{1,1}(\mathbb{R}^{n})$ is included in the Appendix.
@@
Instate all notations in Algorithm 1. Let $p\geq 3$, and let $\alpha_{k}=\frac{p}{k+p}$ for each $k\in\mathbb{N}$. Consider an objective function $f\in\mathscr{F}_{L}^{1,1}(\mathbb{R}^{n})$.
@@
If the objective $f$ is $L$-smooth and satisfies $\inf_{x\in\mathbb{R}^{n}}f(x)>$-$\infty$ ($f$ is possibly nonconvex), then the following holds.
  • For any $\eta\in(0,\frac{2}{3L}]$, the NSA algorithm satisfies
    ${\min_{k\in\{0,1,\cdots,N\}}\|\nabla f(\mathbf{x}_{k})\|^{2}\leq\frac{3(f(\mathbf{x}_{0})-f^{*})}{2\eta(N+1)}}.$
@@
Let $f$ be $L$-smooth and let $\eta\in(0,\frac{2}{3L}]$. Write $\delta_{k}:=f(\mathbf{x}_{k})-f^{*}$ for simplicity. Let $\mathbf{x}_{k},\mathbf{y}_{k}$ be governed by the NSA algorithm. Then for all $k=0,1,2,\cdots$, it holds that
  • $\delta_{k+1}\leq\delta_{k}$, and more specifically $f(\mathbf{x}_{k+1})\leq f(\mathbf{x}_{k})-\frac{2\eta}{3}\|\nabla f(\mathbf{x}_{k})\|^{2}$;
  • $f(\mathbf{x}_{k+1})\leq f(\mathbf{y}_{k})-\frac{2\eta}{3}\|\nabla f(\mathbf{y}_{k})\|^{2}$.
@@
$\delta_{k+1}\leq f(\mathbf{x}_{k+1}^{\prime\prime})-f(\mathbf{x}^{*})\leq f(\mathbf{x}_{k})-\frac{2\eta}{3}\|\nabla f(\mathbf{x}_{k})\|^{2}-f(\mathbf{x}^{*})=\delta_{k}-\frac{2\eta}{3}\|\nabla f(\mathbf{x}_{k})\|^{2},$
which proves the first item. By (2) we know
$f(\mathbf{x}_{k+1})\leq f(\mathbf{x}_{k+1}^{\prime})\leq f(\mathbf{y}_{k})-\frac{2\eta}{3}\|\nabla f(\mathbf{y}_{k})\|^{2},$
which proves the second item. ∎
@@
Proof of the first two items in Theorem 1.
@@
where $\gamma_{k+1}=\frac{\alpha_{k}^{2}}{\eta}$ for all $k=0,1,\cdots$. Thus for any $k\geq 1$, we have
@@
The Advantages of Ensuring Descending. Beyond matching the $o(k^{-2})$ convergence rate of existing methods [doi:10.1137/15M1046095], the NSA method offers the distinct advantage of explicitly ensuring a descent property. This guarantees convergence to a local minimum even for nonconvex objectives, as shown below in Proposition 2.
Proposition 2.
Suppose the objective $f$ is nonconvex and $L$-smooth. If $\eta\in(0,\frac{2}{3L}]$, then it holds that
$\min_{k\in\{0,1,\cdots,N\}}\|\nabla f(\mathbf{x}_{k})\|^{2}\leq\frac{3(f(\mathbf{x}_{0})-f^{*})}{2\eta(N+1)}.$
Proof.
By Lemma 1, we have
$f(\mathbf{x}_{k+1})\leq f(\mathbf{x}_{k})-\frac{2\eta}{3}\|\nabla f(\mathbf{x}_{k})\|^{2},$
By telescoping, the above inequality implies
$\min_{k\in\{0,1,\cdots,N\}}\frac{2\eta}{3}\|\nabla f(\mathbf{x}_{k})\|^{2}\leq\frac{2\eta}{3(N+1)}\sum_{k=0}^{N}\|\nabla f(\mathbf{x}_{k})\|^{2}\leq\frac{f(\mathbf{x}_{0})-f(\mathbf{x}_{N+1})}{N+1}\leq\frac{f(\mathbf{x}_{0})-f^{*}}{N+1}.$